# Removal Attack and Defense on AI-generated Content Latent-based Watermarking

### v2

@@ 1. Introduction
Specifically, an AIGC method, for example a Latent Diffusion Model (LDM), can be viewed as a random source parameterized by a condition. Let us write the generating process as ${\tt L}(\mathbf{c},\mathbf{s})$ where $\mathbf{c}$ is the condition and $\mathbf{s}$ a random starting point sampled from a multivariate Gaussian distribution in a high dimensional (e.g. $d=4\times 64\times 64)$ real vector space. Inversion based AIGC watermarking (Zhao et al., 2024a; Yang et al., 2024b) replaces the original starting point in AIGC with crafted “starting points” drawn from a secret distribution. During detection, given an object, it is first inverted to the starting point and then decides whether the inverted point indeed follows the secret distribution. Unlike traditional watermarking approaches, this method on AIGC does not require minimum distortion of a given host to embed the watermark. Furthermore, by feeding in starting points derived from pseudorandom error correcting (PRC) codes where the codewords are computationally indistinguishable from uniformly sampled binary sequences (Zhao et al., 2024a), it is possible to attain indistinguishable (aka undetectable) watermarked objects.
@@
Removal attacks can be carried out on the victim at different stages of the generation process. It can be on the final generated object in the pixel domain (Zhao et al., 2024a), intermediate latent points (Müller et al., 2024), or all the way back to the starting point in the latent space. In this paper, we consider a removal attack that distorts the starting point in the latent space. Such consideration is reasonable. Firstly, the secret is injected into the starting point and thus it would be insightful to understand its robustness against malicious attacks. Secondly, in practice, the attacker could have more compute resources or other background information that allow for more accurate inversion. For security analysis of a defense mechanism, the threat model should assume an attacker with high capability and able to invert exactly. On the other hand, to analyze the attack effectiveness from an attacker’s perspective, we should also investigate attackers with various levels of capability. Hence, our analysis on the defense mechanism consider attackers who can invert exactly (Theorem 3), and the empirical studies considers inversion both with and without errors (Section 6.2).
We formulate a removal attack scenario in which the attackers have access to multiple copies of watermarked objects (referred as known watermark attack), and compare the attack effectiveness against a navïe whitenoise attacker. Under the scenario, we propose an attack targeting indistinguishable watermark, specifically PRC  (Zhao et al., 2024a) and Gaussian Shading (Yang et al., 2024b). The attack incurs significantly smaller perturbation when compared to whitenoise that attains the same probability of success. Furthermore, the attack is ”stealthy”, in the sense that the attacked starting point remains indistinguishable from the original (to someone who does not know the seed). Under reasonable parameters (Figure 2 (a)), the perturbation can be reduced by a factor of $15\times$. Essentially, this attack attempts to maximize the number of bits flipped in the Hamming space, while incurring small distortion in the real vector space, at the same time ensuring that changes remain indistinguishable. In a certain sense, the significant gain is attained by exploiting the fact that an watermarked object’s location could leak information about its boundary, in particular, the direction of the nearest non-watermarked object. Hence, even if the locations are indistinguishable from random points, effective removal could still be possible.
@@ 3.1.1. False Alarm
$\mathbb{P}[{det}(k,Gauss())=1]\leq\alpha,$
@@ 4.1. Constructions of $\langle smp,{det}\rangle$
Table 1 summarizes constructions by PRC (Zhao et al., 2024a) and Gaussian Shading (Yang et al., 2024b). Both generate binary sequences in step (1), and then together with additional truly random sources, transform it to multivariate Gaussian sample in step (2).
@@
PRC (Zhao et al., 2024a) (1) Generate $b_{1},\dots,b_{d}$, a $d$-bit pseudorandom codeword for the given message. (2) Sample $z_{1},\dots,z_{d}$ independently from $\mathcal{N}(0,1)$ and set $x_{i}=(-1)^{1-b_{i}}|z_{i}|$. (3) During detection, given a latent point, compute the sign of each coefficient and then employ the error-correcting decoder to recover from error.
@@ 6. Performance Analysis
We evaluate the performance of our proposed watermark removal attacks on the Pseudorandom Codeword (PRC) watermarking scheme, utilizing the code implementation from (Christ and Gunn, 2024). The PRC watermarking scheme is parameterized by message length $m$, which is the bit-lenght of messages that can be embedded, and false positive rate (FPR) $f$, which is the probability that the decoder fails to output a message. To convert the scheme to a watermark detector, a message is designated as watermarked and the rest as non-watermarked. Hence, the false alarm, i.e. the probability $p$ that a random latent vector is detected as watermarked is $(1-f)\times 2^{-m}$. To be consistent with the empirical studies in (Zhao et al., 2024a), we set the FPR parameter to be $f=0.01$. Since (1f)(1-f) is close to $1$, we can approximate the false alarm as $2^{-m}$.
@@
We employed Stable Diffusion 2.1 (Rombach et al., 2021) with its default parameters as the latent diffusion model, downloaded from HuggingFace [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1. Images were generated using prompts sourced from the Stable Diffusion Prompt (SDP) dataset [^3] https://huggingface.co/datasets/Gustavosta/Stable-Diffusion-Prompts, using the downloaded model’s default parameters over 50 timesteps. This is consistent with the evaluations performed in prior work (Zhao et al., 2024a; Yang et al., 2024b). In the additional evaluations performed in Appendix C, we did the same but on the Stable Diffusion 1.5444https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 and Stable Diffusion XL [^5] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 models, both also using their default parameters.
@@ 6.1.3. Robustness under Secret Transformation
In this section, we evaluate our proposed attacks and defenses on Stable Diffusion 2.1 (the same model used in the evaluation of (Zhao et al., 2024a)). Further evaluations on Stable Diffusion 1.5 and Stable Diffusion XL are provided in Appendix C. We consider three attacker capabilities (AC) that differ in their inversion accuracy.
@@
We employ the PRC watermark detector implementation by (Zhao et al., 2024a), which is as follows: Given a watermarked image $I$, the detector performs DDIM inversion (conditional on the null prompt) to obtain the approximate latent $s$. Watermark detection is then done on $s$,
@@ 8.1. Digital Watermarking
    Latent space watermarking: Embeds watermarks in the latent space starting point (Zhao et al., 2024a; Yang et al., 2024b; Wen et al., 2023). However, some methods, like Tree-Ring Watermarking (Wen et al., 2023), alter the latent distribution, making them detectable via steganographic attacks (Yang et al., 2024a). Empirical evaluations (Zhao et al., 2024a) also show that Gaussian Shading (Yang et al., 2024b) introduces visually perceptible artifacts in the generated images. To the best of our knowledge, the PRC watermarking scheme (Zhao et al., 2024a) is the only provably undetectable watermarking method and hence, is the focus of this paper.
@@ 8.2. Watermark Removal Attacks
Recently, there has been increasing interest in attacks targeting the removal of watermarks from the outputs of machine learning models (Lukas et al., 2021). These attacks, known as watermark removal attacks, aim to perturb or alter watermarked outputs in a way that renders the embedded watermark undetectable while preserving the usability of the generated content (Pang et al., 2024). For instance, (Zhao et al., 2024b) demonstrated that traditional pixel-level digital watermarking schemes are vulnerable to regeneration attacks, where a specially trained generative AI model creates a semantically similar but unwatermarked version of the image.
To address the limitations of post-hoc pixel-level watermarking, recent research has explored embedding watermarks into the initial latent starting point of the generative process (Yang et al., 2024b; Zhao et al., 2024a; Wen et al., 2023). By integrating the watermark at the latent level, the resulting scheme becomes more robust to common attacks such as noise addition, filtering, and geometric distortions, which typically compromise traditional pixel-level watermarking techniques. To the best of our knowledge, no prior work has specifically addressed the removal of watermarks embedded in the latent space.
@@ Acknowledgements.
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-029).
@@ Appendix A Code Implementation
The code to reproduce our experiments together with instructions is available in the following repository: https://github.com/dezhanglee/watermark_removal. This code largely reuses the code base from the original PRC watermarking paper (Zhao et al., 2024a), which is distributed under the MIT License.

### v3

@@ Abstract.
Watermark, Latent diffusion model, Watermark removal
@@ 1. Introduction
Specifically, an AIGC method, for example a Latent Diffusion Model (LDM), can be viewed as a random source parameterized by a condition. Let us write the generating process as ${\tt L}(\mathbf{c},\mathbf{s})$ where $\mathbf{c}$ is the condition and $\mathbf{s}$ a random starting point sampled from a multivariate Gaussian distribution in a high dimensional (e.g. $d=4\times 64\times 64)$ real vector space. Inversion based AIGC watermarking (Gunn et al., 2024; Yang et al., 2024b) replaces the original starting point in AIGC with crafted “starting points” drawn from a secret distribution. During detection, given an object, it is first inverted to the starting point and then decides whether the inverted point indeed follows the secret distribution. Unlike traditional watermarking approaches, this method on AIGC does not require minimum distortion of a given host to embed the watermark. Furthermore, by feeding in starting points derived from pseudorandom error correcting (PRC) codes where the codewords are computationally indistinguishable from uniformly sampled binary sequences (Christ and Gunn, 2024; Gunn et al., 2024), it is possible to attain indistinguishable (aka undetectable) watermarked objects.
@@
Removal attacks can be carried out on the victim at different stages of the generation process. It can be on the final generated object in the pixel domain (Zhao et al., 2024), intermediate latent points (Müller et al., 2024), or all the way back to the starting point in the latent space. In this paper, we consider a removal attack that distorts the starting point in the latent space. Such consideration is reasonable. Firstly, the secret is injected into the starting point and thus it would be insightful to understand its robustness against malicious attacks. Secondly, in practice, the attacker could have more compute resources or other background information that allow for more accurate inversion. For security analysis of a defense mechanism, the threat model should assume an attacker with high capability and able to invert exactly. On the other hand, to analyze the attack effectiveness from an attacker’s perspective, we should also investigate attackers with various levels of capability. Hence, our analysis on the defense mechanism consider attackers who can invert exactly (Theorem 3), and the empirical studies considers inversion both with and without errors (Section 6.2).
We formulate a removal attack scenario in which the attackers have access to multiple copies of watermarked objects (referred as known watermark attack), and compare the attack effectiveness against a navïe whitenoise attacker. Under the scenario, we propose an attack targeting indistinguishable watermark, specifically PRC  (Gunn et al., 2024) and Gaussian Shading (Yang et al., 2024b). The attack incurs significantly smaller perturbation when compared to whitenoise that attains the same probability of success. Furthermore, the attack is ”stealthy”, in the sense that the attacked starting point remains indistinguishable from the original (to someone who does not know the seed). Under reasonable parameters (Figure 2 (a)), the perturbation can be reduced by a factor of $15\times$. Essentially, this attack attempts to maximize the number of bits flipped in the Hamming space, while incurring small distortion in the real vector space, at the same time ensuring that changes remain indistinguishable. In a certain sense, the significant gain is attained by exploiting the fact that an watermarked object’s location could leak information about its boundary, in particular, the direction of the nearest non-watermarked object. Hence, even if the locations are indistinguishable from random points, effective removal could still be possible.
@@ 3.1.1. False Alarm
$\mathbb{P}[{det}(k,Gauss())=1]\leq\gamma,$
@@ 4.1. Constructions of $\langle smp,{det}\rangle$
Table 1 summarizes constructions by PRC (Gunn et al., 2024) and Gaussian Shading (Yang et al., 2024b). Both generate binary sequences in step (1), and then together with additional truly random sources, transform it to multivariate Gaussian sample in step (2).
@@
PRC (Gunn et al., 2024) (1) Generate $b_{1},\dots,b_{d}$, a $d$-bit pseudorandom codeword for the given message. (2) Sample $z_{1},\dots,z_{d}$ independently from $\mathcal{N}(0,1)$ and set $x_{i}=(-1)^{1-b_{i}}|z_{i}|$. (3) During detection, given a latent point, compute the sign of each coefficient and then employ the error-correcting decoder to recover from error.
@@ 6. Performance Analysis
We evaluate the performance of our proposed watermark removal attacks on the Pseudorandom Codeword (PRC) watermarking scheme, utilizing the code implementation from (Christ and Gunn, 2024). The PRC watermarking scheme is parameterized by message length $m$, which is the bit-lenght of messages that can be embedded, and false positive rate (FPR) $f$, which is the probability that the decoder fails to output a message. To convert the scheme to a watermark detector, a message is designated as watermarked and the rest as non-watermarked. Hence, the false alarm, i.e. the probability $p$ that a random latent vector is detected as watermarked is $(1-f)\times 2^{-m}$. To be consistent with the empirical studies in (Gunn et al., 2024), we set the FPR parameter to be $f=0.01$. Since (1f)(1-f) is close to $1$, we can approximate the false alarm as $2^{-m}$.
@@
We employed Stable Diffusion 2.1 (Rombach et al., 2021) with its default parameters as the latent diffusion model, downloaded from HuggingFace [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1. Images were generated using prompts sourced from the Stable Diffusion Prompt (SDP) dataset [^3] https://huggingface.co/datasets/Gustavosta/Stable-Diffusion-Prompts, using the downloaded model’s default parameters over 50 timesteps. This is consistent with the evaluations performed in prior work (Gunn et al., 2024; Yang et al., 2024b). In the additional evaluations performed in Appendix C, we did the same but on the Stable Diffusion 1.5444https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 and Stable Diffusion XL [^5] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 models, both also using their default parameters.
@@ 6.1.3. Robustness under Secret Transformation
In this section, we evaluate our proposed attacks and defenses on Stable Diffusion 2.1 (the same model used in the evaluation of (Gunn et al., 2024)). Further evaluations on Stable Diffusion 1.5 and Stable Diffusion XL are provided in Appendix C. We consider three attacker capabilities (AC) that differ in their inversion accuracy.
@@
We employ the PRC watermark detector implementation by (Gunn et al., 2024), which is as follows: Given a watermarked image $I$, the detector performs DDIM inversion (conditional on the null prompt) to obtain the approximate latent $s$. Watermark detection is then done on $s$,
@@ 8.1. Digital Watermarking
    Latent space watermarking: Embeds watermarks in the latent space starting point (Gunn et al., 2024; Yang et al., 2024b; Wen et al., 2023). However, some methods, like Tree-Ring Watermarking (Wen et al., 2023), alter the latent distribution, making them detectable via steganographic attacks (Yang et al., 2024a). Empirical evaluations (Gunn et al., 2024) also show that Gaussian Shading (Yang et al., 2024b) introduces visually perceptible artifacts in the generated images. To the best of our knowledge, the PRC watermarking scheme (Gunn et al., 2024) is the only provably undetectable watermarking method and hence, is the focus of this paper.
@@ 8.2. Watermark Removal Attacks
Recently, there has been increasing interest in attacks targeting the removal of watermarks from the outputs of machine learning models (Lukas et al., 2021). These attacks, known as watermark removal attacks, aim to perturb or alter watermarked outputs in a way that renders the embedded watermark undetectable while preserving the usability of the generated content (Pang et al., 2024). For instance, (Zhao et al., 2024) demonstrated that traditional pixel-level digital watermarking schemes are vulnerable to regeneration attacks, where a specially trained generative AI model creates a semantically similar but unwatermarked version of the image.
To address the limitations of post-hoc pixel-level watermarking, recent research has explored embedding watermarks into the initial latent starting point of the generative process (Yang et al., 2024b; Gunn et al., 2024; Wen et al., 2023). By integrating the watermark at the latent level, the resulting scheme becomes more robust to common attacks such as noise addition, filtering, and geometric distortions, which typically compromise traditional pixel-level watermarking techniques. To the best of our knowledge, no prior work has specifically addressed the removal of watermarks embedded in the latent space.
@@ Acknowledgements.
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-029). We also thank the anonymous reviewers for their constructive feedback, and Wenjie Qu for noticing a citation mistake.
@@ Appendix A Code Implementation
The code to reproduce our experiments together with instructions is available in the following repository: https://github.com/dezhanglee/watermark_removal. This code largely reuses the code base from the original PRC watermarking paper (Gunn et al., 2024), which is distributed under the MIT License.